OpenStudy (anonymous):

Find the sum or difference. 1/2-1/14 A 3/7 B 2/7 C 4/7

OpenStudy (sleepyjess):

What is the least common denominator of 2 and 14?

OpenStudy (sleepyjess):

Do you know? If not that is fine. @cutegirlll364

OpenStudy (anonymous):

14

OpenStudy (sleepyjess):

Correct, so we will leave 1/14 the same and multiply 1/2 by 7/7 to get a denominator of 14.

OpenStudy (anonymous):

but non of my answer choices have no denominator with 14

OpenStudy (sleepyjess):

You will have to simplify when we are done

OpenStudy (anonymous):

so it will be 3/7

OpenStudy (sleepyjess):

No, did you multiply \(\dfrac{1}{2}*\dfrac{7}{7}\)

OpenStudy (anonymous):

oh it 2/7

OpenStudy (sleepyjess):

Wait, yes it is \(\dfrac{3}{7}\)

OpenStudy (anonymous):

so I was right about 3/7

OpenStudy (sleepyjess):

Yes, you were correct.